OpenStudy (anonymous):

A gift of $9000 to a city grew to $4000000 in 200 years. At what interest rate compounded annually would this growth occur?

OpenStudy (anonymous):

\[9000(1+r)^{200}=4000000\]solve for \(r\)

OpenStudy (anonymous):

\[(1+r)^{200}=\frac{4000}{9}\] \[1+r=\sqrt[200]{\frac{4000}{9}}\] and now a calculator

OpenStudy (anonymous):

do i dstribute my 9000 to the 1 and the r first

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ohok

OpenStudy (anonymous):

it is \(9000(1+r)^{200}\) you cannot distribute

OpenStudy (anonymous):

ok... subtract my 9000 to both sides right

OpenStudy (anonymous):

divide both side by 9000

OpenStudy (anonymous):

the power of 200 confuses me

OpenStudy (anonymous):

that gives \((1+r)^{200}=\frac{4000}{9}\)

OpenStudy (anonymous):

yes, it should you have to take the 200th root here is the answer http://www.wolframalpha.com/input/?i=%284000%2F9%29^%281%2F200%29

OpenStudy (anonymous):

you get \(1+r=1.0309\) so \(r=.0309\)

OpenStudy (anonymous):

i missed the sqrt part of ur message

OpenStudy (anonymous):

or as a percent about \(3.1\%\)